# Implementation of time-dependent Hartree Fock in real space

### v1

@@ 4.2 Starting point for SCF convergence of HF
Figure 2 shows the convergence characteristics of the SCF procedure for different starting guesses using Broyden density mixing. It is evident from Figure 2 that starting from LDA solution not only makes the convergence faster but is also responsible for any convergence at all as the HF from LCAO didn’t converge within 300 iterations and was not making progress.
Figure 2: Convergence of HF ground-state for thioformaldehyde ($h=0.22$, $R=5$) from different starting points, measured by the relative density error defined in equation 14. First LDA calculation (black) was performed starting from a Linear Combination of Atomic Orbitals (LCAO) and the converged results were used as initial guess for HF calculation (red). The blue curve shows convergence when the HF calculation was performed directly from LCAO.
@@ 4.3 Mixing for SCF convergence of HF
Figure 4: Convergence, as measured by the relative density error defined in equation 14, while mixing states (red and black with different mixing parameters) using linear scheme for the nitrosomethane molecule ($h=0.16$, $R=5$). Blue and magenta curves are linear mixing of potential and density respectively shown as a comparison to states mixing.
@@ 4.4 Convergence of Orbitals with the Real-Space Grid Parameters
To obtain TDHF spectra converged with respect to the grid spacing $h$ and sphere radius $R$, we must converge not only the ground-state density, but also the occupied and unoccupied orbitals’ eigenvalues and wavefunctions. We show the convergence of occupied (Figure 5) and unoccupied (Figure 6) eigenvalues from HF and LDA with respect to the change in the grid spacing of the calculation domain.
The occupied orbitals converge within $0.01$ eV at $0.14$ for both LDA and HF, and have similar convergence behavior. The unoccupied orbitals similarly converge within $0.01$ eV at $0.14$ for LDA, but for HF a finer grid of spacing $0.1$ is required to achieve the same level of convergence. This is not surprising since there is an additional integral to be converged in real space, in HF vs LDA. The occupied orbitals converge monotonically as the spacing is decreased in both LDA and HF. For unoccupied orbitals, this is the case for LDA too but not for HF. In Figure 7 we examine also the convergence of occupied eigenvalues with respect to the sphere radius $R$, which shows that eigenvalues are very well converged by $R=10$. This is as expected since that large radius was chosen to ensure adequate description of the unoccupied orbitals. It is not straightforward to make a meaningful assessment of unoccupied eigenvalues’ convergence with radius, because the state of a given index will continually decrease toward zero as the radius is increased, and converge only to zero. These vacuum states behave similarly to the particle-in-a-box model. Therefore we do not provide a plot of unoccupied states’ convergence with radius. Overall we conclude that convergence of HF unoccupied orbitals is harder in real space but still achievable.
Figure 5: Relative error of the occupied orbitals’ eigenvalues compared to the finest grid ($h=0.09$) at various grid spacings for thioformaldehyde on a domain of union of spheres of radii $5$. Dashed and solid lines represent LDA and HF respectively.
Figure 6: Relative error of the unoccupied orbitals’ eigenvalues compared to the finest grid spacing ($h=0.09$  Å) at various grid spacings for thioformaldehyde on a domain of union of spheres of radii $5$. Dashed and solid lines represent LDA and HF respectively.
Figure 7: Relative error of the occupied orbitals’ eigenvalues compared to the biggest simulation sphere ($R=10$) at various radii for thioformaldehyde molecule with grid spacing $h=0.08$. Dashed and solid lines represent LDA and HF respectively.
@@ 4.5 Benchmarking TDHF
For benchmarking purposes, we’ve chosen to compare our real-space excitation energies from Octopus to those from [50] which uses the GTO code Psi4 [51]. Out of the $17$ molecules from the GTO dataset studied there (from [5]), we excluded formaldehyde, methamine and streptocyanine-c1 from our comparison due to problems with convergence of unoccupied levels in real space. We have compared the excitation energies obtained from these two different codes for IPA and TDA from TDDFT using the BLYP functional and also from TDHF.
The unoccupied orbitals and their symmetry is sensitive to the choice of basis set. The number of virtual orbitals obtainable depends on the number of basis functions in GTOs while in real-space it depends on the number of grid-points. For example, using aug-cc-pvtz basis set on acetaldehyde gives 496 basis functions [50], meaning a total of 496 eigenstates can be calculated. But in real-space we are not limited to that number and finer grids give us a larger number of eigenstates. The extra states obtained in real-space calculation are mostly low-lying Rydberg states. This makes the process of targeting the same excitation for comparison in GTO and real-space based approaches difficult. The indices of the states targeted in [50] are different compared to the corresponding ones from real space. The size of the simulation domain also affects the nature of the calculated orbitals. Increasing the domain size results in introduction of new low-lying vacuum states which may not contribute to the absorption spectrum but nonetheless are important in understanding other kinds of response [16]. Figure 8 a) shows the state targeted for the triplet excitation as obtained using GTOs, which is the fourth state after the lowest unoccupied molecular orbital (LUMO). However, in real-space, the same state index represents a Rydberg orbital as shown in figure 8 b). The state resembling the GTO LUMO+4 state is found in real-space as the $30^{\rm th}$ unoccupied state. With this issue in mind, for the comparison of excitation energies against the data from [50], we have ensured that we are targeting the same excitation from both approaches by making sure the symmetry of the unoccupied orbitals matches from both approaches along with their IPA excitation energies. It is harder to determine the symmetry of the orbitals in real space. The symmetry of GTOs can be used to classify orbital symmetries, typically in the solution process. In real space, the Cartesian grid inherently breaks the system’s symmetry in most cases. In this work, we have manually examined the molecular orbitals to determine their irreducible representations.
@@
Table 2 summarizes the mean error (ME) and the mean absolute error (MAE) of RS excitation energies as compared to the GTO approach from both TDDFT and TDHF for singlets and triplets. In tables 3 and 4, we show all the triplet and singlet excitation energies, respectively, as obtained from TDDFT (TDBLYP) and TDHF, in the IPA and TDA forms, from both the GTO-based Psi4 code and the real-space (RS) Octopus code. The TBE column represents the “Theoretical Best Estimate” from GTO coupled-cluster calculations in [5]. The singlet energies have a lower overall deviation in real-space while GTOs perform better for triplets. This can be attributed to the cancellation of errors between the two integrals in equation 6. The results are also presented in figures 9 and 10 where as error bars we have depicted the signed deviation from the TBE.
We have also calculated the absorption spectrum using TDHF and compared it against the spectrum from TDDFT. The absorption spectra of the nitrosomethane molecule is shown in figure 11. In addition, to show that these spectra can be successfully and systematically converged, we have shown the absorption spectrum for thioformaldehyde at different grid spacings in figure 12.
@@
Table 2: Mean Error (ME) and Mean Absolute Error (MAE) in eV between real-space (RS) Octopus and Gaussian-type orbital (GTO) Psi4 calculations, for excitation energies from various approaches.
@@
Figure 9: Comparison of TDDFT excitation energies of different molecules from GTOs and real-space. The green line represents the points where $\omega^{\rm RS}=\omega^{\rm GTO}.$
@@
Figure 10: Comparison of TDHF excitation energies of different molecules from GTOs and real-space. The green line represents the points where $\omega^{\rm RS}=\omega^{\rm GTO}.$
Figure 11: Comparison of singlet absorption spectra of nitrosomethane molecule from TDDFT using BLYP and TDHF. ($R=10$, $h=0.10$ for both TDDFT and TDHF). Lorentzian of width $0.0136$ eV used for broadening.
Figure 12: Convergence of triplet absorption spectra of thioformaldehyde molecule with respect to grid spacing from TDHF using TDA, with sphere $R=5$. Lorentzian of width $0.0136$ eV used for broadening.
@@ 5 Conclusion
In this work, we have demonstrated a new implementation of linear-response TDHF in the real-space Octopus code, working in the Tamm-Dancoff approximation. For a benchmark set of small molecules, we find good agreement with results from GTOs, though there is a larger discrepancy for TDHF than for the HF eigenvalues or for TD-BLYP calculations. The source of discrepancies is attributed to the use of semilocal PBE pseudopotentials for our calculations; possible differences in degree of convergence of the RS and GTO calculation; and different treatment of Rydberg excitations – which in general are better described in real space than in GTOs where their energies may be overestimated. Although we have only shown results from RHF here, we have also implemented TDHF for UHF and confirmed that the set of excitation obtained from RHF and UHF are identical for these closed-shell molecules. We find that convergence with respect to the real-space domain’s radius and spacing is harder than for DFT, but still achievable.
@@
The calculation of HF ground-state also requires the calculation of exact exchange which is particularly time-consuming. We have analyzed the use of ACE, an approach to speeding up the exact exchange calculation which can also be used for hybrid DFT. Using the existing implementation of ACE in Octopus, we found that for our benchmark set of molecules, the use of ACE sped up the time per SCF iteration by a factor of 7-14. We also note that the little discrepancy between the results from the two basis choices might be because of the unavoidable use of pseudopotentials in real-space. We have also made sure that the results using ACE match the results without it.
This work adds TDHF to the list of features of the Octopus code, allowing us to compare various approaches in the same platform. We also note that with some further developments to the code, one could include hybrid kernels for TDDFT linear-response calculation now, and also enables use of the TDHF bare-exchange integrals in other theories like ensemble DFT.

### v2

@@ 4.2 Starting point for SCF convergence of HF
Figure 2 shows the convergence characteristics of the SCF procedure for different starting guesses using Broyden density mixing. It is evident from Figure 2 that starting from the LDA solution not only makes the convergence faster but is also responsible for any convergence at all, as the HF from LCAO didn’t converge within 300 iterations and was not making progress.
Figure 2: Convergence of HF ground-state for thioformaldehyde ($h=0.22$, $R=5$) from different starting points, measured by the relative density error defined in equation 14. First an LDA calculation (black) was performed starting from a Linear Combination of Atomic Orbitals (LCAO) and the converged results were used as initial guess for HF calculation (red). The blue curve shows convergence when the HF calculation was performed directly from LCAO.
@@ 4.3 Mixing for SCF convergence of HF
Figure 4: Convergence, as measured by the relative density error defined in equation 14, while mixing states (red and black with different mixing parameters) using linear scheme for the nitrosomethane molecule ($h=0.16$, $R=5$). Blue and magenta curves are linear mixing of potential and density, respectively, shown as a comparison to states mixing.
@@ 4.4 Convergence of Orbitals with the Real-Space Grid Parameters
To obtain TDHF spectra converged with respect to the grid spacing $h$ and sphere radius $R$, we must converge not only the ground-state density, but also the occupied and unoccupied orbitals’ eigenvalues and wavefunctions. We show the convergence of occupied and unoccupied eigenvalues in figure 5 from HF and LDA with respect to the change in the grid spacing of the calculation domain.
The occupied orbitals converge within $0.01$ eV at $0.14$ for both LDA and HF, and have similar convergence behavior. The unoccupied orbitals similarly converge within $0.01$ eV at $0.14$ for LDA, but for HF a finer grid of spacing $0.1$ is required to achieve the same level of convergence. This is not surprising since there is an additional integral to be converged in real space, in HF vs LDA. The occupied orbitals converge monotonically as the spacing is decreased in both LDA and HF. For unoccupied orbitals, convergence is monotonic for LDA too, but not for HF. In Figure 6 we examine also the convergence of occupied eigenvalues with respect to the sphere radius $R$, which shows that eigenvalues are very well converged by $R=10$. This is as expected since that large radius was chosen to ensure adequate description of the unoccupied orbitals. It is not straightforward to make a meaningful assessment of unoccupied eigenvalues’ convergence with radius, because the state of a given index will continually decrease toward zero as the radius is increased, and converge only to zero. These vacuum states behave similarly to the particle-in-a-box model. Therefore we do not provide a plot of unoccupied states’ convergence with radius. Overall we conclude that convergence of HF unoccupied orbitals is harder in real space but still achievable.
Refer to captiona) OccupiedRefer to captionb) Unoccupied
Figure 5: Relative error of the occupied and unoccupied orbitals’ eigenvalues compared to the finest grid ($h=0.09$) at various grid spacings for thioformaldehyde on a domain of union of spheres of radii $5$. Dashed and solid lines represent LDA and HF respectively.
Figure 6: Relative error of the occupied orbitals’ eigenvalues compared to the biggest simulation sphere ($R=10$) at various radii for ethylene molecule with grid spacing $h=0.08$. Dashed and solid lines represent LDA and HF respectively.
@@ 4.5 Benchmarking TDHF
For benchmarking purposes, we’ve chosen to compare our real-space excitation energy calculations in Octopus to those from [49] which uses the GTO code Psi4 [50]. Out of the $17$ molecules from the GTO dataset studied there (from [5]), we excluded formaldehyde, methamine and streptocyanine-c1 from our comparison due to problems with convergence of unoccupied levels in real space. We have compared the excitation energies obtained from these two different codes for IPA and TDA from TDHF as well as from TDDFT (using the BLYP functional).
The unoccupied orbitals and their symmetry is sensitive to the choice of basis set. The number of virtual orbitals obtainable depends on the number of basis functions in GTOs while in real-space it depends on the number of grid-points. For example, using aug-cc-pvtz basis set on acetaldehyde gives 496 basis functions [49], meaning a total of 496 eigenstates can be calculated. But in real-space we are not limited to that number and finer grids give us a larger number of eigenstates. The extra states obtained in real-space calculation are mostly low-lying Rydberg states. This makes the process of targeting the same excitation for comparison in GTO and real-space based approaches difficult. The indices of the states targeted in [49] are different from those of the corresponding states in a real space calculation. The size of the simulation domain also affects the nature of the calculated orbitals. Increasing the domain size results in introduction of new low-lying vacuum states which may not contribute to the absorption spectrum but nonetheless are important in understanding other kinds of response [15]. Figure 7a) shows the state targeted for the triplet excitation as obtained using GTOs, which is the fourth state after the lowest unoccupied molecular orbital (LUMO). However, in real-space, the same state index represents a Rydberg orbital as shown in figure 7b). The state resembling the GTO LUMO+4 state is found in real-space as the $30^{\rm th}$ unoccupied state. With this issue in mind, for the comparison of excitation energies against the data from [49], we have ensured that we are targeting the same excitation from both approaches by making sure the symmetry of the unoccupied orbitals matches from both approaches along with their IPA excitation energies. It is harder to determine the symmetry of the orbitals in real space. The symmetry of GTOs can be used to classify orbital symmetries, typically in the solution process. In real space, the Cartesian grid inherently breaks the system’s symmetry in most cases. In this work, we have manually examined the molecular orbitals to determine their irreducible representations.
@@
Table 2 summarizes the mean error (ME) and the mean absolute error (MAE) of RS excitation energies as compared to the GTO approach from both TDDFT and TDHF for singlets and triplets. The MAE for TDHF is 2 times larger than TDDFT for singlets and 5 times for triplets, showing that exact exchange integrals introduce a greater discrepancy. Triplets have a larger MAE than singlets, which we attribute in part to favorable cancellation of errors for the singlet between the two integrals of equation 6, whereas only one term is present for the triplet. MAEs are about an order of magnitude bigger for HF eigenvalues (IPA) than for BLYP, and HF eigenvalues’ MAE is 50-100% of the TDHF MAE. This comparison indicates that the eigenvalues, coming from the long-standing and well-tested HF implementation, rather than the TDHF integrals implemented in this work, are responsible for a large fraction of the discrepancies for TDHF. These discrepancies therefore are considered to stem from the inherent numerical differences between the RS and GTO calculations. In tables 3 and 4, we show all the triplet and singlet excitation energies, respectively, as obtained from TDDFT (TDBLYP) and TDHF, in the IPA and TDA forms, from both the GTO-based Psi4 code and the real-space (RS) Octopus code. The TBE column represents the “Theoretical Best Estimate” from GTO coupled-cluster calculations in [5].
The results are summarized in figures 8 and 9, where the error bars represent the signed deviation from the TBE. We observe that in the case of triplet excitations, TDHF with GTOs is closer to the TBE than TDHF with with real space, whereas for singlet excitations, TDHF with real space is closer to the TBE than TDHF with GTOs. These differing results for TDHF stem from differences in the accuracy of two integrals in equation 6, and cancellation of errors due to basis set and level of theory. By contrast, TDDFT results show similar deviations from the TBE for both GTO and real space.
We have also calculated the absorption spectrum using TDHF and compared it to the spectrum from TDDFT. The absorption spectra of the nitrosomethane molecule is shown in figure 10. In addition, to show that these spectra can be successfully and systematically converged, we have shown the absorption spectrum for thioformaldehyde at different grid spacings in figure 11. We see that convergence of occupied eigenvalues to  0.001 eV and unoccupied eigenvalues to  0.01 eV (figure 5) leads to convergence of singlet excitation energies within  0.1 eV: the integrals in equation 6 are more demanding to converge than the eigenvalues.
@@
Table 2: Mean Error (ME) and Mean Absolute Error (MAE) in eV between real-space (RS) Octopus and Gaussian-type orbital (GTO) Psi4 calculations, for excitation energies from TDDFT (BLYP) and TDHF, in the IPA and TDA approaches.
@@
Figure 8: Comparison of BLYP excitation energies, from the independent particle approximation (IPA) and TDDFT in the Tamm-Dancoff approximation (TDA), as calculated with GTOs [49] and real-space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
@@
Figure 9: Comparison of TDHF excitation energies, from the independent particle approximation (IPA) and the Tamm-Dancoff approximation (TDA), as calculated with GTOs [49] and real-space (RS) for the set of 14 small molecules. The green line represents the line $\omega^{\rm RS}=\omega^{\rm GTO}$, and error bars represent the signed deviation from the “Theoretical Best Estimate” (TBE) [5].
Figure 10: Comparison of absorption spectra of nitrosomethane molecule from TDDFT using BLYP, and TDHF, both using the TDA ($R=10$, $h=0.10$ for both TDDFT and TDHF). Lorentzian of width $0.0136$ eV used for broadening.
Figure 11: Convergence of absorption spectra of thioformaldehyde molecule with respect to grid spacing from TDHF using TDA, with sphere $R=5$. Lorentzian of width $0.0136$ eV used for broadening.
@@ 5 Conclusion
In this work, we have demonstrated a new implementation of linear-response TDHF in the real-space Octopus code, working in the Tamm-Dancoff approximation. For a benchmark set of small molecules, we find good agreement with results from GTOs. The discrepancy between RS and GTOs for TDHF is significantly larger than for TD-BLYP calculations, showing the more sensitive nature of the TDHF integrals. It is also larger for triplets than for singlets. The discrepancy is largely accounted for by discrepancies in the HF eigenvalues, confirming that our TDHF implementation is comparable to the pre-existing HF implementation in Octopus. These discrepancies between RS and GTOs are attributed to the unavoidable use of pseudopotentials in real space; the fact that these pseudopotentials for TDHF are generated with LDA rather than HF; possible differences in degree of convergence of the RS and GTO calculations; and different treatment of Rydberg excitations. We note that in general Rydberg excitations are better described in real space than in GTOs, where their energies may be overestimated, as seen by the negative mean error of RS vs GTOs in many cases (Table 2). Although we have only shown results from RHF here, we have also implemented TDHF for UHF and confirmed that the set of excitation obtained from RHF and UHF are identical for these closed-shell molecules. We find that convergence with respect to the real-space domain’s radius and spacing is harder than for DFT, but still achievable, for both excitation energies and absorption spectra.
@@
The calculation of HF ground-state also requires the calculation of exact exchange which is particularly time-consuming. We have analyzed the use of ACE, an approach to speeding up the exact exchange calculation which can also be used for hybrid DFT. Using the existing implementation of ACE in Octopus, we found that for our benchmark set of molecules, the use of ACE sped up the time per SCF iteration by a factor of 7-14. We have also confirmed that the results using ACE match the results without it.
This work adds TDHF to the list of features of the Octopus code, allowing users to compare various approaches in the same platform. We also note that with some further developments to the code, the ingredients from TDHF can now be used to implement hybrid functionals in TDDFT linear-response calculations [8] and other theories like ensemble DFT [51, 49].